OpenStudy (anonymous):

Sole for x and y e^(2x) + e^(2y) = 10 e^(x+y) = 4 hint: Let p=e^x and q=e^y thanks

OpenStudy (anonymous):

Substituting \(p=e^x\) and \(q=e^y\), you have \[\begin{cases}p^2+q^2=10\\pq=4\end{cases}\] The second equation is due to the fact that \(e^{x+y}=e^xe^y=pq\), in case you were wondering. First, solve for either \(p\) or \(q\) in the second equation. I'll work with \(p\): \[p=\frac{4}{q}\] Next, plug this into the first equation: \[\left(\frac{4}{q}\right)^2+q^2=10\\ \frac{16}{q^2}+q^2=10\\ 16+q^4=10q^2\\ q^4-10q^2+16=0\\ (q^2-8)(q^2-2)=0\\ (q-\sqrt8)(q+\sqrt8)(q-\sqrt2)(q+\sqrt2)=0\] So, \(q=\pm\sqrt8,~\pm\sqrt2\). Now solve for \(p\).